Kohler Corporation reports the following components of stockholders equity on December 31, 2013: 



Common Stock-$10 par value, 100,000 shares authorized, 40,000 shares issued and outstanding.$400,000
Paid-in capital in excess of par value, common stock60,000
Retained Earnings270,000
Total Stockholders' Equity
$730,000

 

In 2014, the following transactions affected its stockholders' equity accounts. 


Jan. 1: Purchased 4,000 shares of its own stock at $20 cash per share. 

Jan. 5: Directors declared a $2 per share cash dividend payable on Feb. 28 to the stockholder of record. 

Feb 28: Paid the dividend declared on January 5. 

July 6: Sold 1,500 of its treasury shares at $24 cash per share. 

Aug. 22: Sold 2,500 of its treasury shares at $17 cash per share. 

Sept. 5: Directors declared a S2 per share cash dividend payable on October 28 to the September 25 stockholders of record.

Oct 28: Paid the dividend declared on September 5. 

Dec. 31: Closed the $388,000 credit balance (from net income) in the Income Summary account to Retained Earnings.


Required: 

.1. Prepare journal entries to record each of these transactions for 2014. 

2. Prepare a statement of retained earnings for the year ended December 31, 2014 

3. Prepare the stockholders' equity section of the company's balance sheet as of December 31, 2014.
Answer #1
1
Date General Journal Debit Credit
1-Jan Treasury stock, Common 80,000 =4000*20
Cash 80,000
5-Jan Retained earnings 72,000 =(40000-4000)*2
Common dividend payable 72,000
28-Feb Common dividend payable 72,000
Cash 72,000
6-Jul Cash 36,000 =1500*24
Treasury stock, Common 30,000 =1500*20
Paid-in capital, Treasury stock 6,000
22-Aug Cash 42,500 =2500*17
Paid-in capital, Treasury stock 6,000
Retained earnings 1,500
Treasury stock, Common 50,000 =2500*20
5-Sep Retained earnings 80,000 =40000*2
Common dividend payable 80,000
28-Oct Common dividend payable 80,000
Cash 80,000
31-Dec Income summary 388,000
Retained earnings 388,000
2
KOHLER CORPORATION
Statement of Retained Earnings
For Year Ended December 31, 2014
Retained earnings, Dec. 31, 2013 270,000
Add: Net income 388,000
658,000
Less: Cash dividends declared (152,000)
Less: Treasury stock reissuances (1,500)
Retained earnings, Dec. 31, 2014 504,500
3
KOHLER CORPORATION
Stockholders' Equity Section of the Balance Sheet
December 31, 2014
Common stock - $10 par value 400,000
Paid-in capital in excess of par value, common stock 60,000
Total contributed capital 460,000
Retained earnings 504,500
Total stockholders' equity 964,500
Answer #2
Date General journal Debit Credit
1-Jan Treasury stock 80,000
cash (4000*20) 80,000
5-Jan Retained earnings 72000
Dividends payable 72,000
(40000-4000)*2
28-Feb Dividends payable 72,000
cash 72,000
6-Jul Cash   (1,500*24) 36000
Treasury stock (1,500*20) 30000
paid in capital in excess of Treasury stock 6000
22-Aug Cash (2500*17) 42500
paid in capital in excess of par -TS 6,000
Retained earnings 1,500
Treasury stock   (2500*20) 50000
5-Sep Retained earnings 80000
Dividends payable (40000*2) 80,000
28-Oct Dividends payable 80,000
cash 80,000
31-Dec Income summary 388,000
Retained earnings 388,000
Statement of retained earnings
opening balance ,jan 1 270,000
add :Net income 388,000
658,000
less:Cash Dividend 152000
        Treasury stock 1,500
Closing balance,Dec 31 504,500
Stockholders Equity section of the Balance Sheet
common stock 400,000
paid in capital in excess of par-CS 60,000
total contributed capital 460,000
Retained earnings 504,500
total Stockholders Equity 964,500